The exercise involving data in this and subsequent sections were designed to be solved using Excel.

The following estimated regression equation is based on 30 observations.

y= 18.1 +4x1 - 2.4x2 +7.7x3+2.6x4

The values of SST and SSR are 1807 and 1759, respectively.

a. Compute r2 (to 3 decimals).

b. Computer ra2 (to 3 decimals).
Answer #1

a) r2 = SSR/SST = 1759 / 1807 = 0.973

b) ra2 = 1 - (1 - 0.973)*(30 - 1) / (30 - 4 - 1) = 0.969